# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v3

@@ 9.9.2 Model-Based Generation of a Sequence of Contact Points and Forces/Torques
As shown in Figure 10, we use a simple process to explain the generation of a sequence of contact points and forces/torques. Assuming the motion trajectory of the object is already obtained, the target wrench can be calculated based on the object’s mass and inertia. Contact points can be optimized based on object geometry, object material, and end-effector geometry etc, using appropriate metrics (Ferrari et al. (1992)). Alternatively, they can be generated using a knowledge-based approach (Stansfield (1991)). Subsequently, a coordinate system is established at the centroid of the object. Based on the location of the contact points, the target wrench is converted into fingertip force. Finally, the fingertip force is converted into the forces/torques required by the actuators through hand jacobian (Okamura et al. (2000)). The pipeline contact points and forces/torques mentioned above are derived sequentially. In practical applications, both can be optimized simultaneously (Xu et al. (2024)).
Rolling and sliding at the contact points can hinder task execution. Tasks like pushing and pulling don’t need strong stability between the end-effector and the object. However, for tasks like grasping, stability is crucial and the system must keep stable contact even when disturbed. Form-closure and force-closure are common approaches. Form-closure depends on the end-effector’s design. Force-closure involves using force control to counteract external forces, like gravity. Current grasping methods mainly increase grip force to boost friction and achieve force-closure. However, to achieve precise wrench control, it is essential to understand the coefficient of friction represented by the object’s material. Tactile-based slip prediction can also enhance stability (Veiga et al. (2015)), but it requires high real-time performance.
@@
The process described above represents a simple pipeline. However, it still faces significant challenges in contact-rich tasks. For model-based contact-rich tasks, the challenges include: (1) the explosion of contact modes, (2) stiff and non-smooth contact dynamics, and (3) the non-convexity of the planning problem (Pang et al. (2023)). Reinforcement learning (RL) has shown success in contact-rich manipulation tasks through efficient sampling and averaging of contact modes. Pang et al. (2023) develop a mathematical model to demonstrate the theoretical equivalence between random smoothing and analytic smoothing. This work achieves results comparable to RL and requires significantly less computation.
@@ 9.9.3 Learning-Based Generation of a Sequence of Contact Points and Forces/Torques
One major challenge in data collection for dexterous manipulation lies in gathering data from multi-fingered end-effectors. Teleoperation requires a real-robot system, which is not portable and cannot achieve in-the-wild data collection. Therefore, current mainstream research focuses on directly tracking human hand motions during manipulation without controlling the robot.
@@
Current learning methods each have their strengths and weaknesses. Therefore, learning approaches for dexterous manipulation should integrate different methods. For example, diffusion policy can assist reinforcement learning in addressing high-dimensional action spaces issue, while reinforcement learning can help diffusion policy overcome issues with suboptimal and negative data. Additionally, the learning models should consider both inputs and outputs. The factors necessary for achieving dexterous manipulation are summarized in the ”Model-Based Generation of a Sequence of Contact Points and Forces/Torques”. ManiFoundation(Xu et al. (2024)) generates contact point and force heatmaps based on task goals, object material, object geometry and end-effector geometry. However, ManiFoundation lacks sensors such as tactile for real-time servoing of the manipulated object’s motion.
@@ 9.10 How to establish a benchmark?
Current research on foundation models for manipulation focuses on various tasks, including interaction, hierarchical tasks, perception, detecting pre- and post-conditions, policy, and manipulation data generation. Therefore, a benchmark for foundation models for manipulation should include a comprehensive framework with diverse tasks. This framework should test individual tasks and tasks that involve connecting different modules. Building a benchmark should include existing datasets. However, since different simulators have unique physics engines and renderers, the benchmark should include various simulators and separately test all existing datasets in different simulators.
@@
As for the metric for assessing general manipulation. The current evaluation standards mainly focus on success rates. However, in real-world applications, user preferences should also be considered. For instance, the system’s real-time performance is important. Most algorithms focus on building the generalization of skills. They often overlook the amount of data and speed required for RFMs to learn a new skill. Therefore, evaluation metric should also include the learning ability of RFMs.

### v4

@@ 9.9.2 Model-Based Generation of a Sequence of Contact Points and Forces/Torques
As shown in Figure 10, we use a simple process to explain the generation of a sequence of contact points and forces/torques. Assuming the motion trajectory of the object is already obtained, the target wrench can be calculated based on the object’s mass and inertia. Contact points can be optimized based on object geometry, object material, and end-effector geometry etc, using appropriate metrics (Ferrari et al. (1992)). Alternatively, they can be generated using a knowledge-based approach (Stansfield (1991)). Subsequently, a coordinate system is established at the centroid of the object. Based on the location of the contact points, the target wrench is converted into fingertip force. Finally, the force of the fingertip is converted into the forces/torques required by the actuators through hand jacobian (Okamura et al. (2000)).
The pipeline contact points and forces/torques mentioned above are derived sequentially. In practical applications, both can be optimized simultaneously (Xu et al. (2024)). However, fingertips do not always make contact with the object’s surface through optimization. This results in a non-smooth contact dynamic nature and creates challenges for planning in contact-rich manipulation (Pang (2023)). As shown in Fig. 11, $q^{a}$ and $\mu$ represent the actual and commanded positions of the actuated ball, while $q^{u}$ represents an unactuated box that slides along a rail with sufficient damping. Our goal is to push the box by the ball to transition from the current configuration $q^{u}$ to the goal configuration $q^{u}_{goal}$. Suppose $q^{u}_{+}$ is the steady-state position of the box after the ball is commanded to position $\mu$. We expect that $q^{u}_{+}$ and $q^{u}_{goal}$ are sufficiently close. Typically, $q^{u}_{+}=f(q^{u},q^{a},\mu)$ represents the system’s dynamic, which can be obtained through system identification (Suh et al. (2023)).
As shown in Fig. 11, a gradient exists only when the ball and the box are in contact, which presents challenges for optimization-based problem-solving methods. Current approaches for smoothing contact dynamics are mainly divided into analytic smoothing, first-order randomized smoothing, and zeroth-order randomized smoothing (Pang et al. (2023)). However, it is often necessary to consider frictional contact constraints, which can make the optimization problem non-convex. To address this issue, current methods primarily rely on Anitescu’s convex relaxation (Todorov et al. (2012); Pang et al. (2023); Todorov (2014); Anitescu (2006); Castro et al. (2022)). At the same time, the optimization should avoid robot joint limits, penetration between the hand and object, and self-penetration as much as possible (Pang and Tedrake (2021); Xu et al. (2024)). Although smoothed dynamics on trajectory optimization, such as iMPC (Suh et al. (2022) and LQR (Shirai et al. (2024)) are considered effective, these methods are only suitable when the goal is close to the initial configuration. When the goal is farther from the initial configuration, a more global strategy is required.
As shown in Fig. 11, there are two initial guesses: 1) the ball is on the left of the box, and 2) the ball is on the right of the box. Different initial guesses lead to different dynamic model. Choosing the wrong initial guess can prevent the optimization problem finding the optimal solution. To address this issue, the system dynamics need to be divided into contact modes, and a global strategy is required (Pang et al. (2023)). Fig. 11 includes four contact modes: left contact, left no contact, right contact, and right no contact. However, for a system with $n$ rigid bodies, there can be up to $C^{2}_{n}$ contact pairs. Each contact pair may have multiple contact modes, such as sticking, separation, and sliding, which can lead to an explosion of contact modes. Although Huang et al. (2021) considers kinematically feasible contact modes and reduces the number to $O(n^{d^{2}/2+2.5d})$(where d is the number of degrees of freedom and n is the number of contact points), the reduced number of contact modes is still too large to handle.
To address the above issues, current model-based techniques are mainly divided into contact-explicit and contact-implicit (Jiang et al. (2024b)). The contact-explicit approach in dexterous in-hand manipulation involves explicitly representing contacts, such as their locations, modes, and reaction forces (Cheng et al. (2023)). These representations form contact sequences, which are obtained through methods like enumerating (Aceituno-Cabezas and Rodriguez (2020); Hogan and Rodriguez (2020)), searching (Chen et al. (2021a); Zhu et al. (2023a); Cruciani et al. (2018)), sampling (Cheng et al. (2023)), or demonstration (Khadivar and Billard (2023)). However, these methods suffer from poor scalability (Pang et al. (2023)) and contacts easily fall into the local optimum (Jiang et al. (2024b)). The contact-implicit approach avoids the complexity of explicit contact representation. It uses methods like relaxed complementary constraints (Kim et al. (2023); Le Cleac’h et al. (2024)), smooth surrogate models (Pang et al. (2023); Önol et al. (2019)), or direct control sampling (Howell et al. (2022)). However, the smoothing process introduces force-at-a-distance effects and sacrifices physical fidelity (Pang et al. (2023)). Although Jiang et al. (2024b) compensates for discrepancies between actual and planned contact modes using tactile feedback, this method assumes quasi-static contact model and struggles to handle highly dynamic actions, such as spinning a pen between fingers.
@@
Figure 11: Non-smooth dynamic system.
@@ 9.9.3 Learning-Based Generation of a Sequence of Contact Points and Forces/Torques
One major challenge in data collection for dexterous manipulation lies in gathering data from multi-fingered end-effectors. Although model-based hard code method (Zhu et al. (2024)) can collect data on dexterous manipulation, they still require data analysis such as mutual information (Hejna et al. (2025)) and entropy (Zhu et al. (2024)) to assess the quality of the data. Additionally, for multi-scenario and multi-task data collection, teleoperation methods are less dependent on algorithm performance compared to model-based hard code methods. However online teleoperation requires a real-robot system, which is not portable and cannot achieve in-the-wild data collection. Therefore, current mainstream research focuses on directly tracking human hand motions during manipulation without controlling the real-robot.
@@
Current learning methods each have their strengths and weaknesses (Zhang et al. (2024a)). Therefore, learning approaches for dexterous manipulation should integrate different methods. For example, diffusion policy can assist reinforcement learning in addressing high-dimensional action spaces issue, while reinforcement learning can help diffusion policy overcome issues with suboptimal and negative data. Additionally, the learning models should consider both inputs and outputs. The factors necessary for achieving dexterous manipulation are summarized in the ”Model-Based Generation of a Sequence of Contact Points and Forces/Torques”. ManiFoundation(Xu et al. (2024)) generates contact point and force heatmaps based on task goals, object material, object geometry and end-effector geometry. However, ManiFoundation lacks sensors such as tactile for real-time servoing of the manipulated object’s motion.
@@ 9.10 How to establish a benchmark?
Current research on foundation models for manipulation focuses on various tasks, including interaction, hierarchical tasks, perception, detecting pre- and post-conditions, policy, and manipulation data generation. Therefore, a benchmark for foundation models for manipulation should include a comprehensive framework with diverse tasks. This framework should test individual tasks and tasks that involve connecting different modules. Since different simulators have unique physics engines and renderers, the benchmark should include standardized simulators and datasets.
@@
As for the metric for assessing general manipulation. The current evaluation standards mainly focus on success rates. However, in real-world applications, other metrics should also be considered. For instance, the system’s real-time performance is important. Most algorithms focus on building the generalization of skills. They often overlook the amount of data and speed required for RFMs to learn a new skill. Therefore, evaluation metric should also include the learning ability of RFMs.